OpenStudy (timaashorty):

Are you stalking my profile ? Don't get confused. I'm talking to YOU. YES You. Who else ? You're the only one reading this!

OpenStudy (anonymous):

Tima you know how to do this pellet btw?

OpenStudy (shamil98):

i like the universe

OpenStudy (anonymous):

All the questions you've been asking?

OpenStudy (shamil98):

you have a vertical stretch of 1/4 and a vertical shift of-2 units (down 2 )

OpenStudy (shamil98):

and the graph is pointed downwards

OpenStudy (timaashorty):

Yess i do, or i could simply search them ^.^ But I'm trying to get Shams ss Up.

OpenStudy (shamil98):

its either B or D, i need to check the inverse.

OpenStudy (anonymous):

get mine up

OpenStudy (anonymous):

I need 70

OpenStudy (timaashorty):

Well LEts see you'll have the quickest answers then ;P lol

OpenStudy (timaashorty):

Lmao i put universe again i just noticed

OpenStudy (shamil98):

y=-4x^2-2 x = -4y ^2 - 2 x + 2 = -4y^2 √(x+2/-4) = y

OpenStudy (anonymous):

It's universe is a parabola

OpenStudy (anonymous):

It's not b sham :d

OpenStudy (shamil98):

yeah the graph is pointing to the right its d.

OpenStudy (anonymous):

i dont think so but meh to lazy to check

OpenStudy (anonymous):

Prob A

OpenStudy (shamil98):

no, y =- 4x^2 - 2 the coeffecient is negative, it points downwards

OpenStudy (shamil98):

and the vertical shift is -2, not + 2 :P

OpenStudy (anonymous):

Oh I didn't see the negative sign

OpenStudy (anonymous):

-4

OpenStudy (shamil98):

yep so D

OpenStudy (timaashorty):

Guys make sure. if i see get anything wrong. you'll both get buttwhipping ;c

OpenStudy (timaashorty):

i still have a bad grade in this =c

OpenStudy (shamil98):

math is e z

OpenStudy (anonymous):

ye it's prob d

OpenStudy (anonymous):

sham wants the d

OpenStudy (timaashorty):

I'll go with D. Then tell me goodluck

OpenStudy (anonymous):

http://www.youtube.com/watch?v=svrKVImuzhk

OpenStudy (timaashorty):

Lool Batman you do too

OpenStudy (timaashorty):

uh oh nin came ;o

OpenStudy (nincompoop):

your math is way advanced too advanced it is stephen hawking-level talking about universe and multi-verse

OpenStudy (anonymous):

pfft just cause he has high ss, doesn't mean shat.

OpenStudy (timaashorty):

i think he wants it too

OpenStudy (timaashorty):

Ikr Ninn ;P I'm in here levels thank you all.